OpenStudy (anonymous):

(medal + fan and a cookie :) ) 3. Solve for h. A=1/2bh a. h= 2A/b b. 2A= b/h c. h= 1/2Ab d. h= 2Ab explain pleaseee :)

OpenStudy (anonymous):

so the answer is A? O_O

OpenStudy (anonymous):

ans is optn (C)

OpenStudy (anonymous):

no it was by mistake .. correct ans is optn (c)

OpenStudy (anonymous):

can you explain so I could get it and know.

OpenStudy (anonymous):

\[A=\frac{ 1 }{ 2bh }\] \[2bh=\frac{ 1 }{ A } \] \[h=\frac{ 1 }{ 2Ab }\]

OpenStudy (anonymous):

could I ask one more question?

OpenStudy (anonymous):

ya sure ..

OpenStudy (anonymous):

2. Solve for y, then find the value of y when given x = -2 A. -12 B. 7/4 C. 19/4 D. 24

OpenStudy (anonymous):

where is the equation of y ?

OpenStudy (anonymous):

OH oops. the equation is 6x= 7-4y

OpenStudy (anonymous):

4y=7-6x \[y=\frac{ 7-6x }{ 4 }\] now putting x=-2.. we get y=19/4